# Imprints of interaction processes in the globular cluster system of NGC 3640

### v1

@@ 3.3 Colour distribution
In Figure 9 we show the histogram of the colour distribution for the GCs attributed to NGC 3640, as well as the background correction. Since the background sources show a homogeneous distribution across the colour range, we consider the background correction negligible in terms of fitting the subpopulations.
Figure 10 shows the density probability functions of the colour distribution for the GCs attributed to NGC 3640, as well as the inner ($r_{gal}<75^{\prime\prime}$) and outer ($r_{gal}>75^{\prime\prime}$) regions. These smoothed distributions use a Gaussian kernel, with a bandwidth of $0.5$.
@@
Both regions are statistically bimodal according to this fit. The means of the subpopulations in the inner region are redder than when considering the entire sample, in particular for the red subpopulation. The outer region presents bluer means for both subpopulations than the entire sample, and is more likely to be contaminated by GCs connected to NGC 3641.
@@
Figure 10: Smoothed olour distributions for the full sample of GC candidates attributed to NGC 3640 (top). In the middle panel, the inner region (left), and in the bottom, the outer region of NGC 3640. Blue solid curves and red dashed curves represent the best fit from GMM for bimodal Gaussian distributions for each sample.
@@ 3.4 Luminosity function
The luminosity function of globular cluster systems is widely used as a distance estimator due to the universality of the absolute magnitude of its turn-over. The SBF distance to NGC 3640 is $26.9$ Mpc, which results in a turn-over magnitude of $V_{0}=24.75$. Using the transformations from Faifer et al. (2011), and a mean $(g-i)$ value of 1, we can transform to approximately $i_{0}=24.1$, which is well within our completeness range, allowing us to estimate our own measurement of the distance to compare with the literature value.
In Figure 11, we show the luminosity function for GC candidates associated with NGC 3640, after applying a completeness and background correction. Fitting a Gaussian function to our distribution results in a turn-over value of $i_{0}=23.74\pm 0.02$, which is within error from the SBF distance. The dispersion of the Gaussian is $\sigma_{L}=0.68\pm 0.02$. In Harris et al. (2014), the GCLF is proposed to have an intrinsic width of $\sigma_{L}=0.52-0.55\,\textrm{dex}$. The difference between these two values may be due to contamination from NGC 3641.
Due to this slight difference, we choose to use the turn-over value from SBF to estimate the percentage of GCs within our completeness range in relation to the total population, which in this case is $74\%$. Combining this result with the number obtained from the integration of the radial distribution, we can estimate the total number of GCs for NGC 3640, $N_{GC}\approx 350$.
@@ 4.1 Radial distribution
In Figure 12 we show the radial distribution for the GCs within $75^{\prime\prime}$ of the centre of NGC 3641, with a Hubble-Reynolds law fit up to the imposed limit. If we extrapolate this profile to the same level we used for NGC 3640, we obtain an estimate for the population of NGC 3641 up to the completeness level of $\approx 100$ GCs.
@@ 4.2 Colour distribution
In Figure 13 we show the colour distribution for GCs associated with NGC 3641. Analogously to the previous section, we use GMM and in this case obtain a negative kurtosis ($-0.384$) but the value of D is significantly smaller than 2 ($1.74$), indicating an unimodal distribution is a better fit. The mean for the distribution is $1.05$, with $\sigma=0.16$. In the bimodal fit, the fraction of red GCs is $0.84$, with a mean of $1.09$, showing most of the GCs surrounding NGC 3641 can be considered red. This is reasonable considering we are only analysing objects really close to the galaxy, where red GCs are expected to be dominant, and also because if NGC 3640 is actually stripping its GCs, it would have stripped mainly blue ones which are more external.
@@ 4.3 Luminosity function
In Figure 14 we show the luminosity function of GC candidates associated to NGC 3641. We perform a similar analysis as with NGC 3640, performing the fit of a Gaussian function both using the turn-over from the literature (in this case, the same one as for NGC 3640), and then fitting our own value. We find an almost exact match, although our luminosity function is noisy, with a dispersion larger than usual ($0.9$), and showing bumps towards the brighter end which may be due to contamination from NGC 3640.
An integration of the luminosity function results in a total population of $\approx 135$ GC.
Figure 14: Luminosity function for the GC candidates associated to NGC 3641. Dashed green lines show the background distribution, black solid bars indicate the result from subtracting it from the raw distribution (green bars). The dot-dashed curve shows the Gaussian fit.
@@ 5 Spatial distribution and tidal features
The panels of Figure 15 show the contours of the spatial distribution of the GC candidates located across the field containing both galaxies, separated by colour, placing the limit in $(g^{\prime}-i^{\prime})=1$, drawn on top of residual images. These were obtained modelling both galaxies using the ELLIPSE task in iraf, and subtracting it from the image.
@@
The red GCs are more concentrated towards the region between both galaxies, and their density maps reveal a bridge that connects them. In combination with the overdensity in the colour distribution of NGC 3640 that matches the population of GCs associated to NGC 3641, this seems to indicate that NGC 3640 is accreting GCs from its neighbour, and this process is ongoing.
@@ 6.1 GCs as tracers of substructures
D’Abrusco et al. (2022) study statistically significant structures in the spatial distributions of the GCS of a variety of Fornax galaxies, concluding they are relics of accreted GC systems that are overlaid on the smooth distribution of the GCs belonging to the host galaxy. Following their analysis, we show in Figure 17 the colour distribution for the GCs that correspond spatially to the shells detected in the galaxy, identified in Figure 16. In shell F, located in the direction of NGC 3641, we find most GCs have a similar color, around $0.9$, bluer than those associated to NGC 3641. In substructures A and E, blue GCs dominate, while the region covered by substructure B/C/D has a wider color distribution, although it is also dominated by blue objects, including GCs with colors bluer than $0.7$.
@@ 6.2 Paucity of GCs in the inner region
The Hubble-Reynolds profile fit to the projected radial distribution of the GCS of NGC 3640 has a core radius of $r_{0}=1.16\,\textrm{arcmin}=9.07\,kpc$. Considering NGC 3640 has a stellar mass of $log{M_{\star}}=11.2$ (Karachentsev et al., 2015) we can place it in the context of the $r_{0}$ vs. $M_{\star}$ relation studied in Caso et al. (2024) for a sample of early-type galaxies in HST/ACS archival date. Given that Gemini data is not comparable in terms of depth with HST/ACS, even after a thorough completeness analysis, this number might be slightly enlarged in comparison to those in the sample in Caso et al. (2024). However, it is clear NGC 3640 has a larger core radius than most galaxies in the sample, especially for those with similar masses. This implies the GCS of NGC 3640 is flatter up to a larger galactocentric radius, which can also be seen by simple visual inspection of the projected spatial distribution in Figure 16, where the paucity of GCs near the centre of the galaxy is apparent.
@@ 6.3 The bridge?
The possibility of NGC 3640 and NGC 3641 interacting is not discussed in previous work, mainly due to the observations focusing on the central regions NGC 3640. In terms of integrated light, we do not detect a significant overdensity that can be interpreted as a stream between the two, but it could potentially be fainter than our magnitude limit. Since GCs have been shown to trace stellar streams (Blom et al., 2014; Napolitano et al., 2022, e.g.), it is reasonable to interpret the overdensity detected between the two galaxies as a hint that there might be a stream between them. Previous works in other galaxies have also found potential bridges made of GCs as tracers of interactions, such as Bassino et al. (2006) connecting NGC 1387 and NGC 1399, and Wehner et al. (2008) with NGC 3311 and NGC 3309. We see that the colours of the GCs spatially situated between both galaxies are mainly intermediate, closer to the peak of the population associated with NGC 3641, perhaps due to being members of its GCS that are being accreted by the more massive galaxy. As is originally mentioned in P88, the compactness of NGC 3641 implies a high angular momentum interaction. Although they refrain from interpreting anything more conclusive, they point out that the dust lane in NGC 3640 is oriented in the direction of its neighbour, which is encouraging.
@@ Acknowledgements
This research was supported in part by Perimeter Institute for Theoretical Physics. Research at Perimeter Institute is supported by the Government of Canada through the Department of Innovation, Science and Economic Development and by the Province of Ontario through the Ministry of Research and Innovation. This work was funded with grants from Consejo Nacional de Investigaciones Científicas y Técnicas de la República Argentina, Agencia Nacional de Promoción Cientf́ica y Tecnológica, and Universidad Nacional de La Plata (Argentina). This work made use of AstroPy (Astropy Collaboration et al., 2013, 2018, 2022), MatPlotLib (Hunter, 2007), NumPy (Harris et al., 2020) Pandas (McKinney, 2010), PhotUtils (Bradley et al., 2023), SciKit (Pedregosa et al., 2011) and SciPy (Virtanen et al., 2020).

### v2

@@ 3.3 Colour distribution
In Figure 9 we show the histogram of the colour distribution for the GCs attributed to NGC 3640, as well as the background correction. This correction is estimated by scaling the amount of sources in the background region to the area considered for the GCS. Since the background sources show a homogeneous distribution across the colour range, we consider the background correction negligible in terms of fitting the subpopulations.
Figure 10 shows the density probability functions of the colour distribution for the GCs attributed to NGC 3640, as well as the inner ($r_{gal}<1.25\,$arcmin) and outer ($r_{gal}>1.25\,$arcmin) regions. The distance to separate the regions was estimated by identifying an abrupt drop in the distribution of red GCs at around this value, which could be an indication of a change in the dominating physical process. Slight variations were considered and found to have no significant impact.. These smoothed distributions use a Gaussian kernel, with a bandwidth of $0.5$.
@@
Both regions are statistically bimodal according to this fit. Both subpopulations present redder means in the inner region in comparison to their means across the whole field, with the effect being more noticeable for red GCs. The outer region presents bluer means for both subpopulations than the entire sample, and is more likely to be contaminated by GCs connected to NGC 3641.
@@
Figure 10: In orange curves, smoothed colour distributions for the full sample of GC candidates attributed to NGC 3640 (top). In the middle panel, the inner region (left), and in the bottom, the outer region of NGC 3640. Blue solid curves and red dashed curves represent the best fit from GMM for bimodal Gaussian distributions for each sample.
@@ 3.4 Luminosity function
The luminosity function of globular cluster systems is widely used as a distance estimator due to the universality of the absolute magnitude of its turn-over (e.g. Villegas et al., 2010; Rejkuba, 2012), with a value of $M_{V}=-7.4$. The SBF distance to NGC 3640 is $26.9$ Mpc ($m-M=32.15$), which results in a turn-over magnitude of $V_{0}=24.75$. Using the transformations from Faifer et al. (2011) shown in Equation 3, and a mean $(g-i)$ value of 1, we can transform to approximately $i_{0}=24.1$ mag, which is well within our completeness range, allowing us to estimate our own measurement of the distance to compare with the literature value.
$V_{0}=g_{0}-0.34(\pm 0.02)(g-i)_{0}-0.03(\pm 0.02)$ (3)
In Figure 11, we show the luminosity function for GC candidates associated with NGC 3640. It is derived in the $i^{\prime}$ band since the photometric uncertainties were slightly lower than in the other two. This distribution was corrected using the completeness functions shown in Section 2.3, accounting for radial variations. A background correction was also applied, considering a scaling to the area analysed.. Fitting a Gaussian function to our distribution results in a turn-over value of $i_{0}=23.9\pm 0.2$ mag, where the error was obtained through bootstrapping. The distance modulus for this galaxy according to the SBF distance quoted before is $32.1\pm 0.1$ mag, which results in a turn-over magnitude of $i_{0}=24.1\pm 0.1$ mag. The dispersion of the Gaussian is $\sigma_{i}=0.67\pm 0.2$ mag. In Harris et al. (2014), the GCLF is proposed to have an intrinsic width of $\sigma_{L}=1-1.25\,\textrm{mag}$ for a galaxy of this stellar mass. The lack of complete angular coverage in the GMOS field and the contamination from NGC 3641 might be causing this difference.
We use the turn-over value from SBF and the dispersion of the Gaussian function fit to our distribution using this fixed value, $\sigma_{L}=0.81$ mag (slightly larger than the one obtained in our fit) to estimate the percentage of GCs within our completeness range in relation to the total population, which in this case is $74\%$. Combining this result with the number obtained from the integration of the radial distribution, we can estimate the total number of GCs for NGC 3640, $N_{GC}=362\pm 9$. This is made possible by the three passbands having the same depth at the turn-over magnitude.
Considering the stellar mass of NGC 3640, we can calculate the parameter T as defined by Zepf & Ashman (1993), which is the number of GCs per stellar galaxy mass normalized to $10^{9}M_{\odot}$. We obtain $T=2.3\pm 0.1$. Comparing to Figure 5 in De Bórtoli et al. (2022), which compares $T$ values to the stellar mass of the host galaxy, this is within the expected values for a galaxy of this mass.
@@ 4.1 Radial distribution
In Figure 12 we show the radial distribution for the GCs within 1.251.251.25\,1.25arcmin of the centre of NGC 3641 corrected by completeness and background, with a Hubble-Reynolds law fit up to the imposed limit, resulting in:
$n(r)[N\rm{arcmin}^{-2}]=2.11\left(1+\left(\frac{r[\rm{arcmin}]}{0.69}\right)^{2}\right)^{-2.99}$ (4)
If we extrapolate this profile to the same level we used for NGC 3640, we obtain an estimate for the population of NGC 3641 up to the completeness level of $101\pm 25$ GCs.
@@ 4.2 Colour distribution
In Figure 13 we show the colour distribution for GCs associated with NGC 3641. Analogously to the previous section, we use GMM and in this case obtain a negative kurtosis ($-0.384$) but the value of D is significantly smaller than 2 ($1.74$), indicating an unimodal distribution is a better fit. The mean for the distribution is $(g^{\prime}-i^{\prime})_{0}=1.05$ mag, with $\sigma=0.16$ mag. In the bimodal fit, the fraction of red GCs is $0.84$, with a mean of $1.09$, showing most of the GCs surrounding NGC 3641 can be considered red. This is reasonable considering we are only analysing objects really close to the galaxy, where red GCs are expected to be dominant, and also because if NGC 3640 is actually stripping its GCs, it would have stripped mainly blue ones which are more external.
@@ 4.3 Luminosity function
In Figure 14 we show the luminosity function of GC candidates associated to NGC 3641, corrected by completeness and background estimation. We perform a similar analysis as with NGC 3640, performing the fit of a Gaussian function both using the turn-over from the literature (in this case, the same one as for NGC 3640), and then fitting our own value. We find an almost exact match, $i^{\prime}_{0}=24.11\,\pm 0.3$, although our luminosity function is noisy, with a dispersion slightly smaller than usual ($0.9$ mag) compared to the intrinsic width mentioned above from Harris et al. (2014), and showing bumps towards the brighter end which may be due to contamination from NGC 3640.
In this case, the GCLF indicates that our coverage up to our completeness limit is of $82\%$. Considering the number of GCs estimated from the radial distribution, the mean value corresponding to the SBF distance and the dispersion estimated from our distribution, we calculate a total population of $120\pm 25$ GC.
Figure 14: Luminosity function for the GC candidates associated to NGC 3641. Dashed grey lines show the background distribution, green solid bars indicate the result from subtracting it from the raw distribution (white bars with grey solid lines). The dot-dashed curve shows the Gaussian fit.
@@ 5 Spatial distribution and tidal features
The panels of Figure 15 show the contours of the spatial distribution of the GC candidates located across the field containing both galaxies, separated by colour, placing the limit in $(g^{\prime}-i^{\prime})=1$ mag, drawn on top of residual images. These were obtained modelling both galaxies using the ELLIPSE task in iraf, and subtracting it from the image. For the contours we used a 2D kernel density estimate using a Gaussian kernel, using the sklearn package (Pedregosa et al., 2011).
@@
Whereas blue GCs present overdensities surrounding the galaxies, and seem to be missing from the region closest to NGC 3641, the red GCs around NGC 3641 show an asymmetric distribution, with an overdensity towards NGC 3640 that can be interpreted as a bridge that connects them. This seems to indicate that NGC 3640 is accreting GCs from its neighbour, and this process is ongoing.
@@ 6.1 GCs as tracers of substructures
D’Abrusco et al. (2022) study statistically significant structures in the spatial distributions of the GCS of a variety of Fornax galaxies, concluding they are relics of accreted GC systems that are overlaid on the smooth distribution of the GCs belonging to the host galaxy. Following their analysis, we show in Figure 17 the colour distribution for the GCs that correspond spatially to the shells detected in the galaxy, identified in Figure 16. In shell F, located in the direction of NGC 3641, we find most GCs have a similar color, around $(g^{\prime}-i^{\prime})_{0}=0.9$ mag, bluer than those associated to NGC 3641, where the mean colour was found to be $(g^{\prime}-i^{\prime})_{0}=1.05$ mag. In substructures A and E, blue GCs dominate, while the region covered by substructure B/C/D has a wider color distribution, although it is also dominated by blue objects, including GCs with colors bluer than $(g^{\prime}-i^{\prime})_{0}=0.7$ mag.
@@ 6.2 Paucity of GCs in the inner region
Figure 18: Core radius of the GCS vs. stellar mass of the galaxy for the sample in Caso et al. (2024) in blue circles. A black diamond indicates the position of NGC 3640.
The Hubble-Reynolds profile fit to the projected radial distribution of the GCS of NGC 3640 has a core radius of $r_{0}=1.16\,\textrm{arcmin}=9.07\,kpc$. Considering a stellar mass of $log{M_{\star}}=11.2$ we can place it in the context of the $r_{0}$ vs. $M_{\star}$ relation studied in Caso et al. (2024) for a sample of early-type galaxies in HST/ACS archival date (see Figure 18). Given that Gemini data is not comparable in terms of depth with HST/ACS, even after a thorough completeness analysis, this number might be slightly enlarged in comparison to those in the sample in Caso et al. (2024), and GC detection is typically less efficient in the inner regions of galaxies. However, it is clear NGC 3640 has a larger core radius than most galaxies in the sample, especially for those with similar masses. This implies the GCS of NGC 3640 is flatter up to a larger galactocentric radius, which can also be seen by simple visual inspection of the projected spatial distribution in Figure 16, where the paucity of GCs near the centre of the galaxy is apparent.
@@ 6.3 The bridge?
The possibility of NGC 3640 and NGC 3641 interacting is not thoroughly discussed in previous work, mainly due to the observations focusing on the central regions NGC 3640. In terms of integrated light, we do not detect a significant overdensity that can be interpreted as a stream between the two, but it could potentially be fainter than our magnitude limit. Since GCs have been shown to trace stellar streams (e.g Blom et al., 2014; Napolitano et al., 2022), it is reasonable to interpret the overdensity detected between the two galaxies as a hint that there might be a stream between them. Previous works in other galaxies have also found potential bridges made of GCs as tracers of interactions, such as Bassino et al. (2006) connecting NGC 1387 and NGC 1399, and Wehner et al. (2008) with NGC 3311 and NGC 3309. We see that the colours of the GCs spatially situated between both galaxies are mainly intermediate, closer to the peak of the population associated with NGC 3641, perhaps due to being members of its GCS that are being accreted by the more massive galaxy. As is originally mentioned in P88, the compactness of NGC 3641 might imply a high angular momentum interaction. Although they refrain from interpreting anything more conclusive, they point out that the dust lane in NGC 3640 is oriented in the direction of its neighbour, which is encouraging.
@@ Acknowledgements
We thank the referee for their careful reading of the paper, and for their valuable comments and suggestions. This research was supported in part by Perimeter Institute for Theoretical Physics. Research at Perimeter Institute is supported by the Government of Canada through the Department of Innovation, Science and Economic Development and by the Province of Ontario through the Ministry of Research and Innovation. This work was funded with grants from Consejo Nacional de Investigaciones Científicas y Técnicas de la República Argentina, Agencia Nacional de Promoción Cientf́ica y Tecnológica, and Universidad Nacional de La Plata (Argentina). This work made use of AstroPy (Astropy Collaboration et al., 2013, 2018, 2022), MatPlotLib (Hunter, 2007), NumPy (Harris et al., 2020) Pandas (McKinney, 2010), PhotUtils (Bradley et al., 2023), SciKit (Pedregosa et al., 2011) and SciPy (Virtanen et al., 2020).